### v4

@@ Abstract
Zero-knowledge proof (ZKP) systems have surged attention and held a fundamental role in contemporary cryptography. Zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) protocols dominate the ZKP usage, implemented through arithmetic circuit programming paradigm. However, underconstrained or overconstrained circuits may lead to bugs. The former refers to circuits that lack the necessary constraints, resulting in unexpected solutions and causing the verifier to accept a bogus witness, and the latter refers to circuits that are constrained excessively, resulting in lacking necessary solutions and causing the verifier to accept no witness. This paper introduces a novel approach for pinpointing two distinct types of bugs in ZKP circuits. The method involves encoding the arithmetic circuit constraints to polynomial equation systems and solving them over finite fields by the computer algebra system. The classification of verification results is refined, greatly enhancing the expressive power of the system. A tool, AC4, is proposed to represent the implementation of the method. Experiments show that AC4 demonstrates a increase in the checked ratio, showing a 29% improvement over Picus, a checker for Circom circuits, and a 10% improvement over halo2-analyzer, a checker for halo2 circuits. Within a solvable range, the checking time has also exhibited noticeable improvement, demonstrating a magnitude increase compared to previous efforts.
Keywords:
zero-knowledge proofscircuit constraintsunderconstrainedoverconstrainedalgebraic computation checker
@@ 1 Introduction
Zero-knowledge proof (ZKP) systems have long occupied a fundamental position in contemporary cryptography, ever since their inception [27]. They are significantly utilized in security-sensitive applications, such as smart contracts [53], blockchain confidential transactions [53, 28], and digital currency [60], given their capacity to bolster security and privacy. ZKP systems’ potency lies in their ability to prove anything that an interactive proof system can, but without revealing any knowledge [33, 10]. Therefore, these systems have assumed the role of a potent tool enabling trust and authentication while safeguarding sensitive information.
@@
Developers frequently choose the zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) [12] protocol for its efficiency in setting up and proving within a zero-knowledge proof (ZKP) system. Additionally, they adopt arithmetic circuits [48] as the programming paradigm due to their simplicity in representing computational problems. Both Circom and halo2 are built upon this protocol and programming paradigm. To illustrate this, consider the process of declaring a boolean value in circuit programs. In high-level languages, a boolean value $b$ can be declared directly. However, in circuit languages, a constraint must be explicitly defined to limit its possible values. For instance, the constraint $b\times(b-1)=0$ ensures that $b$ can only be 0 or 1. This transformation cannot be automatically performed by existing comilers or libraries, necessitating that developers express computations using polynomial equations.
@@
The limitation of polynomial constraint systems gives rise to two problems. The first problem is the underconstrained problem [46], which occurs when multiple outputs can fulfill the same system of equations with identical input values. This situation arises when the system of polynomial equations does not yield a mathematical function on the finite field, rendering the circuit underconstrained. As a result, malicious users could generate fake witnesses, deceive the verifier, and potentially steal the cryptocurrency [4, 19]. The second one is the overconstrained problem [46], which occurs when no outputs can satisfy the system of polynomial equations or the constraint system is inconsistent for a given input. This issue can result in a non-functional circuit. Consequently, this may lead to overconstrained bugs that cause ZK-verifier servers to reject numerous legitimate requests. Such rejections could potentially trigger Distributed Denial of Service (DDoS) problems. Furthermore, it is evident that underconstrained circuits can produce more severe consequences than their overconstrained counterparts.
@@
To this end, we conducted an investigation into the constraints on Circom or halo2 circuit programs. By reconsidering the formalization paradigm, we modeled underconstrained and overconstrained problems as algebraic inquiries. Specifically, we examined whether a system of polynomial equations over a finite field has multiple solutions. To support our motivation, we analyzed the official library of circom circuit templates provided by iden3 and collected halo2 circuit templates from . Our analysis revealed that nearly 70% of these templates are linear, and efficient algorithms exist for solving linear systems of equations. Consequently, we categorized constraint problems into diverse types based on their maximal degree. We then employed a unified arithmetic solution framework to address each type separately.
@@
  • We constructed datasets specifically tailored for Circom and halo2 circuits. By comparing the checked ratio and checking speed of AC4 against other tools, we observed substantial improvements in both metrics.
@@ 2 Background
Zk-SNARKs. Non-interactive ZKP protocols, such as zk-SNARKs [11], allow verifiers to be assured of a statement’s truth without gaining any knowledge beyond its validity [27]. Blum et al. demonstrated that computational zero-knowledge can be achieved without interaction by sharing a common random string between the prover and the verifier [15, 58]. The non-interactive protocol conveys the proof through a message from the prover to the verifier, ensuring efficiency. zk-SNARKs reign supreme in ZKP systems due to lightning-fast verification of even large programs, using tiny proofs (mere hundreds of bytes) within milliseconds.
@@
Definition 1 (Arithmetic circuit [55])
@@
Circuit Languages. The growing field of ZKPs has led to the development of specialized circuit languages designed to create and manage the underlying circuits needed for these proofs. Two notable examples are the Circom [9] and halo2 [59] frameworks.
Circom is a domain-specific language that leverages SNARK technology and employs arithmetic circuits as its programming paradigm. Programmers define circuits using specific operators: ======= = = for equality checks, <<--< - - or >-->- - > for signal assignment, and <==absent<==< = = or ==>==>= = > for simultaneous assignment. These circuits conform to the rank-1 constraint system (R1CS) format. The R1CS format is expressed as $A\times B+C=0$, where $A$, $B$, and $C$ are linear combinations of signals. After compiling the circuit into R1CS format and computing a witness using the input, the snarkjs tool can be employed to generate and verify a ZKP for the specified input.
Halo2 provides a more flexible and scalable method for constructing ZKPs. It employs PLONKish arithmetization [26], which supports recursive proof composition, thereby enhancing scalability. In halo2, circuits are built using polynomials and lookups. This construction allows for more efficient proof generation and verification.
@@
where $n$ is the degree of the constraint, $a_{i,j},b_{i},c\in F_{p}$ are the coefficient scalars, and $x_{i}\in F_{p}$ are the variables.
@@
An arithmetic circuit with constraint with constraints can be defined as a tuple $\langle C(\mathcal{X}),E\rangle$, where $C(\mathcal{X})$ is an arithmetic circuit on variable set $\mathcal{X}$. $\mathcal{X}$ is the union of two disjoint sets $K$ and $U$, where $K$ represents known input variables, and $U=T\cup O$ represents unknown variables including the non-constant variables including temporal variables $T$ and output variables $O$. $E$ includes all the constraints of $C(\mathcal{X})$.
Computer algebra. Computer algebra [16] revolves around the manipulation of algebraic expressions and various mathematical entities. This field has spurred the creation of computer algebra systems capable of autonomously tackling a broad spectrum of theoretical and practical mathematical challenges. A contemporary computer algebra system boasts a diverse array of functionalities, encompassing Gröbner bases, cylindrical algebraic decomposition, lattice basis reduction, linear system solving, arbitrary and high-precision arithmetic, interval arithmetic, linear and nonlinear optimization, Fourier transforms, Diophantine equation solving, among others. These capabilities are indispensable for both theoretical investigations and practical applications.
Gröbner basis. A Gröbner basis is a set of multivariate polynomials that has desirable algorithmic properties. Every set of polynomials can be transformed into a Gröbner basis. This process generalizes three familiar techniques: Gaussian elimination for solving linear systems of equations, the Euclidean algorithm for computing the greatest common divisor of two univariate polynomials, and the Simplex algorithm for linear programming. Algorithms on Gröbner basis are widely used in mathematical and verification tools. The computation of a Gröbner basis can be performed using Buchberger’s algorithm [54], while Faugère F5 algorithm [23] is a significant improvement over the former. It introduces the concept of signatures to predict and eliminate unnecessary computations, which can potentially increase the efficiency of Gröbner basis computation.
@@ 3.1 Constraint Problems
Overconstrained bugs. As described in the work by Labs et al. [35], overconstrained circuits are unable to uphold the completeness of ZKP. A circuit program is deemed overconstrained when its corresponding polynomial constraints are overconstrained, indicating that the circuit cannot produce any valid outputs regardless of the inputs provided. This inconsistency in the constraint system for a given input characterizes the state of being overconstrained. Despite their deceptive simplicity in concept and the misconception that they are less critical than other types of attacks, denial of service (DoS) attacks are often underappreciated regarding their security implications. Notably, DoS attacks require neither high skill nor deep knowledge,, contributing to their low cost for malicious attackers as highlighted by Zcash [38].
@@ 3.3 Problem Reduction
Based on the argument above, completely checking the constraint status of circuits can be achieved by solving polynomials. More concretely, the constraint problem corresponds to figuring out the number of solutions for polynomials.
@@ 4.1 AC4 Framework
Fig. 1 illustrates the overall workflow of our proposed tool, $\text{AC}^{4}$, which receives a Circom program file as the input, and compiles the program in debug mode into an R1CS format file along with a sym file. To retrieve the variables and constraints from the compiled files, a task-specific parser is implemented to handle those compiled files, focusing on unknown variables within constraints. After that, the extracted constraints are then simplified into equations. A categorizing mechanism is leveraged in Section 3.2 to determine the specific category of the circuit according to its simplified equation system. Based on the category result of circuits, three types of solvers is employed to address them respectively.
@@
In the context of a precisely pure linear circuit, the system of linear equations can be represented in matrix form. While it may seem natural to rely on rank to determine the number of solutions, it is more important to prioritize verifying the uniqueness of the output variable set. Therefore, solving the system demands a suitable method for linear equations. Specifically, in $\text{AC}^{4}$, the Gauss-Jordan method[51] is utilized to obtain the solution of the circuits as well as to ascertain the uniqueness of the set of output variables.
@@
For higher-order circuits, we solve the coefficients and verify if the circuit will be underconstrained with specific inputs. Then, we use $\text{AC}^{4}$ to directly solve the systems of polynomial equations with Gröebner basis[18] and check the uniqueness of the output variables.
@@
Based on the classification of circuits mentioned above, it is evident that this circuit is not linear. To analyze this, we begin by treating the variable $inp$ as a coefficient and proceed to solve for the undetermined coefficients of each equation. This process yields the solutions [$[\{inp=0\},\{inp=1\}]$. When substituting $\{inp=0\}$ into $E$, the resulting equations are $E=\{0=0,out_{1}\times(inp-1)=0,success\times(success-1)=0\}$. Notably, $out_{0}$ can take on any value over the finite field, indicating that the circuit is underconstrained.
@@ 4.2 Adaptive Checking with Relaxation and Partial Results
[…]
@@ 4.3 Checking Circuits: Category-Based Algorithms
From above, we divide the overall circuit verification problem space into three types: precisely linear, $k$-coefficient linear, and higher-order separately. In this section, we introduce those algorithms in detail.
Precisely linear. The majority of circuits are precisely linear, so they can be represented by a coefficient matrix $A$ and a constant vector $b$. While it is commonly thought to use the rank of the equation matrix to determine the number of solutions, the solver’s objective is only to verify the uniqueness of output variables. In contrast, rank-reducing methods impose restrictions on the uniqueness of all variables, including temporary variables. Using the Gauss-Jordan method to get the rank of the corresponding matrix is analogous to getting the solution of the system of equations. Consequently, there is no significant difference in the time complexity. The pesudo-code is given in Algorithm
@@
1:procedure CheckLinearCircuit($\langle C(\mathcal{X}),E\rangle$)
2:     $A,b\leftarrow\textsc{LinearEquationsToMatrix}(E,U)$triangleright $A$ is the coefficient matrix and $b$ is the constant matrix
@@
6:     else if CheckUniqueness(sols, $O$then
@@
Higher-Order. A heuristic approach will be used to identify input variables that may cause the circuit to become underconstrained (shown in Algorithm3), aiming to improve accuracy and efficiency. The coefficients of the unknown variables are treated as equations and their solutions will be stored in a frequency table for counting the occurrences of each solution. Then, the solver sorts the solutions based on their frequencies and iteratively substitutes them into the original system of equations to obtain a new system, which is then solved to check if the circuit is underconstrained. If the input variable causing the circuit to become underconstrained cannot be identified, the system of equations will be directly solved with Gröbner basis to verify the constrained situation of the circuit. The pseudo-code of checking polynomial circuits is given in Algorithm2.
@@
procedure CheckMultiSolutionsByCoeffs($K,U,E$)
     $UCSolCountTable\leftarrow EmptyHashTable$
     for all $eq\in E$ do $\triangleright$ iterate all the equations
         $sols\leftarrow\textsc{SolveUndeterminedCoefficientEq}(eq,K,U)$
         if $sols\neq\emptyset$ then $\triangleright$ If there exists solution
              for all $sol\in sols$ do
                  $UCSolCountTable[sol]\leftarrow UCSolCountTable[sol]+1$
              end for
         end if
     end for
     return $UCSolCountTable$
end procedure
Algorithm 3 Algorithm for checking higher-order circuits
procedure CheckHigherCircuit($\langle C(\mathcal{X}),E\rangle$)
     do the same thing to undetermined coefficients and check underconstrained
     $sols\leftarrow\textsc{SolveHigherSystemWith$F_{5}$}(C)$
     if IsEmpty(sols) then
         return algebraically overconstrained
     else if CheckUniqueness(sols, $O$) then
         return algebraically exact-constrained
     else if $\textsc{Size}(sols)>1$ then
         return precisely underconstrained
     else
         return unknown
     end if
end procedure
Although in nature, these equations are considered linear due to their higher-order terms containing one or more known variables, Hence, they are treated as linear circuits and represented in matrix form to enhance verification efficiency. For these circuits, we still use the previous heuristic approach given in Algorithm 2, to determine if there are special inputs that can cause the underconstrained condition, aiming to enhance efficiency. Afterwards, a $k$-coefficient linear circuit can be succinctly represented by a coefficient matrix $A$ and a constant vector $b$, and checked as a linear case. The pseudo-code of checking $k$-coefficient circuits is given in Algorithm4.

### v5

@@ Abstract
Zero-knowledge proof (ZKP) systems have surged attention and held a fundamental role in contemporary cryptography. Zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) protocols dominate the ZKP usage, implemented through arithmetic circuit programming paradigm. However, underconstrained or overconstrained circuits may lead to bugs. The former refers to circuits that lack the necessary constraints, resulting in unexpected solutions and causing the verifier to accept a bogus witness, and the latter refers to circuits that are constrained excessively, resulting in lacking necessary solutions and causing the verifier to accept no witness. This paper introduces a novel approach for pinpointing two distinct types of bugs in ZKP circuits. The method involves encoding the arithmetic circuit constraints to polynomial equation systems and solving them over finite fields by the computer algebra system. The classification of verification results is refined, greatly enhancing the expressive power of the system. A tool, AC4, is proposed to represent the implementation of the method. Experiments show that AC4 demonstrates a increase in the solved rate, showing a 29% improvement over Picus and CIVER, and a slight improvement over halo2-analyzer, a checker for halo2 circuits. Within a solvable range, the checking time has also exhibited noticeable improvement, demonstrating a magnitude increase compared to previous efforts.
@@ 1 Introduction
Zero-knowledge proof (ZKP) systems have long occupied a fundamental position in contemporary cryptography, ever since their inception [abc10]. They are significantly utilized in security-sensitive applications, such as smart contracts [abc98], blockchain confidential transactions [abc73], and digital currency [zcash], given their capacity to bolster security and privacy. ZKP systems’ potency lies in their ability to prove anything that an interactive proof system can, but without revealing any knowledge [10.1007/3-540-48184-2_4, 10.1007/0-387-34799-2_4]. Therefore, these systems have assumed the role of a potent tool enabling trust and authentication while safeguarding sensitive information.
@@
Developers frequently choose the zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) [184425] protocol for its efficiency in setting up and proving within a zero-knowledge proof (ZKP) system. Additionally, they adopt arithmetic circuits [algebraicComplexityTheory] as the programming paradigm due to their simplicity in representing computational problems. Both Circom and halo2 are built upon this protocol and programming paradigm. To illustrate this, consider the process of declaring a boolean value in circuit programs. In high-level languages, a boolean value $b$ can be declared directly. However, in circuit languages, a constraint must be explicitly defined to limit its possible values. For instance, the constraint $b\times(b-1)=0$ ensures that $b$ can only be 0 or 1. This transformation cannot be automatically performed by existing compilers or libraries, necessitating that developers express computations using polynomial equations.
@@
The limitation of polynomial constraint systems gives rise to two problems. The first problem is the underconstrained problem [10.1145/3591282], which occurs when multiple outputs can fulfill the same system of equations with identical input values. This situation arises when the system of polynomial equations does not yield a mathematical function on the finite field, rendering the circuit underconstrained. As a result, malicious users could generate fake witnesses, deceive the verifier, and potentially steal the cryptocurrency [DisclosureRecentVulnerabilities, cashTornadoCashGot2019]. The second one is the overconstrained problem [10.1145/3591282], which occurs when no outputs can satisfy the system of polynomial equations or the constraint system is inconsistent for a given input. This issue can result in a non-functional circuit. Consequently, this may lead to overconstrained bugs that cause ZK-verifier servers to reject numerous legitimate requests. Such rejections could potentially trigger Distributed Denial of Service (DDoS) problems. Furthermore, it is evident that underconstrained circuits can produce more severe consequences than their overconstrained counterparts.
@@
To this end, we conducted an investigation into the constraints on Circom or halo2 circuit programs. By reconsidering the formalization paradigm, we modeled underconstrained and overconstrained problems as algebraic inquiries. Specifically, we examined whether a system of polynomial equations over a finite field has multiple solutions. To support our motivation, we analyzed the official library of circom circuit templates provided by iden3 and collected halo2 circuit templates from GitHub Projects. Our analysis revealed that nearly 70% of these templates are linear, and efficient algorithms exist for solving linear systems of equations. Consequently, we categorized constraint problems into diverse types based on their maximal degree. We then employed a unified arithmetic solution framework to address each type separately.
@@
  • We constructed datasets specifically tailored for Circom and halo2 circuits. By comparing the solved rate and checking speed of AC4 against other tools, we observed substantial improvements in both metrics.
@@ 2 Background
Zk-SNARKs. Non-interactive ZKP protocols, such as zk-SNARKs [6956581], allow verifiers to be assured of a statement’s truth without gaining any knowledge beyond its validity [abc10]. Blum et al. demonstrated that computational zero-knowledge can be achieved without interaction by sharing a common random string between the prover and the verifier [10.1145/62212.62222, 560484]. The non-interactive protocol conveys the proof through a message from the prover to the verifier, ensuring efficiency. zk-SNARKs reign supreme in ZKP systems due to lightning-fast verification of even large programs, using tiny proofs (mere hundreds of bytes) within milliseconds.
@@
Definition 1(Arithmetic circuit [pmlr-v49-volkovich16])
@@
Circuit Languages. The growing field of ZKPs has led to the development of specialized circuit languages designed to create and manage the underlying circuits needed for these proofs. Two notable examples are the Circom [Circom] and halo2 [opensourcebook:halo2book] frameworks.
Circom is a domain-specific language that leverages SNARK technology and employs arithmetic circuits as its programming paradigm. Programmers define circuits using specific operators: ====== for equality checks, <<-- or >--> for signal assignment, and <==<== or ==>==> for simultaneous assignment. These circuits conform to the rank-1 constraint system (R1CS) format. The R1CS format is expressed as $A\times B+C=0$, where $A$, $B$, and $C$ are linear combinations of signals. After compiling the circuit into R1CS format and computing a witness using the input, the snarkjs tool can be employed to generate and verify a ZKP for the specified input.
Halo2 provides a more flexible and scalable method for constructing ZKPs. It employs PLONKish arithmetization [plonk], which supports recursive proof composition, thereby enhancing scalability. In halo2, circuits are built using polynomials and lookups. This construction allows for more efficient proof generation and verification.
@@
where $n$ is number of variables, $a_{i,j},b_{i},c\in F_{p}$ are the coefficient scalars, and $x_{i}\in F_{p}$ are the variables.
Remark 1
Since any higher-degree ($\geq 3$) polynomial equation can be reduced to a system of quadratic equations by introducing auxiliary variables, we restrict attention to constraints of degree at most $2$. For example, the constraint $x^{3}+c=0$ can be rewritten as $xy+c=0$ and $x^{2}-y=0$.
@@
An arithmetic circuit with constraint can be defined as a tuple $\langle C(\mathcal{X}),E\rangle$, where $C(\mathcal{X})$ is an arithmetic circuit on variable set $\mathcal{X}$. $\mathcal{X}$ is the union of two disjoint sets $K$ and $U$, where $K$ represents known input variables, and $U=T\cup O$ represents unknown variables including the non-constant variables including temporal variables $T$ and output variables $O$. $E$ includes all the constraints of $C(\mathcal{X})$.
Computer algebra. Computer algebra [DBLP:conf/cascon/BrightKG19] revolves around the manipulation of algebraic expressions and various mathematical entities. This field has spurred the creation of computer algebra systems capable of autonomously tackling a broad spectrum of theoretical and practical mathematical challenges. A contemporary computer algebra system boasts a diverse array of functionalities, encompassing Gröbner bases, cylindrical algebraic decomposition, lattice basis reduction, linear system solving, arbitrary and high-precision arithmetic, interval arithmetic, linear and nonlinear optimization, Fourier transforms, Diophantine equation solving, among others. These capabilities are indispensable for both theoretical investigations and practical applications.
Gröbner basis. A Gröbner basis is a collection of multivariate polynomials possessing favorable algorithmic characteristics. Any set of polynomials can be converted into a Gröbner basis. This transformation encompasses three well-known methods: Gaussian elimination for solving systems of linear equations, the Euclidean algorithm for finding the greatest common divisor of two univariate polynomials, and the Simplex algorithm for linear programming. Algorithms on Gröbner basis are widely used in mathematical and verification tools. The computation of a Gröbner basis can be performed using Buchberger’s algorithm [definition:GrobnerBasis], while Faugère F5 algorithm [f5algorithm] is a significant improvement over the former. It introduces the concept of signatures to predict and eliminate unnecessary computations, which can potentially increase the efficiency of Gröbner basis computation.
@@ 3.1 Constraint Problems
Overconstrained bugs. As described in the work by Labs et al. [labs2023SecurityConcernsZeroKnowledge], overconstrained circuits are unable to uphold the completeness of ZKP. A circuit program is deemed overconstrained when its corresponding polynomial constraints are overconstrained, indicating that the circuit cannot produce any valid outputs regardless of the inputs provided. This inconsistency in the constraint system for a given input characterizes the state of being overconstrained. Despite their deceptive simplicity in concept and the misconception that they are less critical than other types of attacks, denial of service (DoS) attacks are often underappreciated regarding their security implications. Notably, DoS attacks require neither high skill nor deep knowledge,, contributing to their low cost for malicious attackers as highlighted by Zcash [ZcashDenialOfServiceAttack].
@@ 3.3 Problem Reduction
Based on the argument above, the complete checking of the constraint status of the circuits can be achieved by solving polynomials. More concretely, the constraint problem corresponds to figuring out the number of solutions for polynomials.
@@ 4.1 AC4 Framework
Fig. 1 illustrates the overall workflow of our proposed tool, $\text{AC}^{4}$, which receives a circuit as the input, and compiles the program in debug mode into an R1CS format file along with a sym file. To retrieve the variables and constraints from the compiled files, a task-specific parser is implemented to handle those compiled files, focusing on unknown variables within constraints. After that, the extracted constraints are then simplified into equations. A categorizing mechanism is leveraged in Section 3.2 to determine the specific category of the circuit according to its simplified equation system. Based on the category result of circuits, three types of solvers is employed to address them respectively.
@@
In the context of a precisely pure linear circuit, the system of linear equations can be represented in matrix form. While it may seem natural to rely on rank to determine the number of solutions, it is more important to prioritize verifying the uniqueness of the output variable set. Therefore, solving the system demands a suitable method for linear equations. Specifically, in $\text{AC}^{4}$, the Gauss-Jordan method[siahaan2023study] is utilized to obtain the solution of the circuits as well as to ascertain the uniqueness of the set of output variables.
@@
For higher-order circuits, we solve the coefficients and verify if the circuit will be underconstrained with specific inputs. Then, we use $\text{AC}^{4}$ to directly solve the systems of polynomial equations with Gröebner basis[BUCHBERGER2006475] and check the uniqueness of the output variables.
@@
Based on the classification of circuits mentioned above, it is evident that this circuit is not linear. To analyze this, we begin by treating the variable inpinp as a coefficient and proceed to solve for the undetermined coefficients of each equation. This process yields the solutions $\{inp=0\},\{inp=1\}]$. When substituting $\{inp=0\}$ into $E$, the resulting equations are $E=\{0=0,out_{1}\times(inp-1)=0,success\times(success-1)=0\}$. Notably, $out_{0}$ can take on any value over the finite field, indicating that the circuit is underconstrained.
@@ 4.2 Adaptive Checking with Relaxation and Partial Results
In certain cases, the input coefficients in circuit equations can be difficult to determine as zero or non-zero. To simplify, AC4 treats all combinations of input and constant values as non-zero, allowing for a more straightforward algebraic result. This means the coefficient can be reduced in the field, thus yields a workable solution, even if not fully precise. Under this assumption, the overconstrained and exact-constrained results may not align with the actual situation of the circuits. We name the former as algebraic results. In contrast, by precisely, we mean that the result are directed obtained without the above assumption. Hence, as depicted in Fig. 3, AC4 generates five practical categories of results: algebraic overconstrained, precisely overconstrained, algebraic exact-constrained, precisely exact-constrained, and precisely underconstrained.
@@ 4.3 Checking Circuits: Category-Based Algorithms
From above, we divide the overall circuit verification problem space into three types: precisely linear, $k$-coefficient linear and higher-order separately. In this section, we introduce those algorithms in detail.
Precisely linear. The majority of circuits are precisely linear, so they can be represented by a coefficient matrix $A$ and a constant vector $b$. While it is commonly thought to use the rank of the equation matrix to determine the number of solutions, the solver’s objective is only to verify the uniqueness of output variables. In contrast, rank-reducing methods impose restrictions on the uniqueness of all variables, including temporary variables. Using the Gauss-Jordan method to get the rank of the corresponding matrix is analogous to getting the solution of the system of equations. Consequently, there is no significant difference in the time complexity. The pesudo-code is given in Algorithm 1.
@@
2:  $b\leftarrow\textsc{LinearEquationsToMatrix}(E,U) \triangleright AA$ is the coefficient matrix and $b$ is the constant matrix
3:  $sols\leftarrow\textsc{GaussJordanSolve(A, b)}$
@@
6:  else if CheckUniqueness(sols, $O$) then
@@
$K$-coefficient linear. Some circuits are inherently nonlinear but can be treated as linear and represented in matrix form to improve verification efficiency. These equations are considered linear because their higher-order terms contain one or more known variables. For such circuits, we employ the heuristic approach to identify special inputs that may cause underconstrained conditions (shown in Algorithm 2), thereby improving efficiency. Afterwards, a $k$-coefficient linear circuit can be succinctly represented by a coefficient matrix $A$ and a constant vector $b$, and checked as a linear case. The pseudo-code of checking $k$-coefficient circuits is given in Algorithm 3.
@@
1:procedure GetUndeterminedCoeffSols($K,U,E$)
2:  $UCSolCountTable\leftarrow EmptyHashTable$
3:  for all $eq\in E$ do $\triangleright$ iterate all the equations
4:   $sols\leftarrow\textsc{SolveUndeterminedCoefficientEq}(eq,K,U)$
5:   if $sols\neq\emptyset$ then $\triangleright$ If there exists solution
6:     for all $sol\in sols$ do
7:      $UCSolCountTable[sol]\leftarrow UCSolCountTable[sol]+1$
8:     end for
9:   end if
10:  end for
11:  return UCSolCountTableUCSolCountTable
12:end procedure
@@
Higher-Order. For those higher-order circuits which cannot be treated as linear, the previous heuristic approach(Algorithm 2) will still be used to identify input variables that may cause the circuit to become underconstrained. The coefficients of the unknown variables are treated as equations and their solutions will be stored in a frequency table for counting the occurrences of each solution. Then, the solver sorts the solutions based on their frequencies and iteratively substitutes them into the original system of equations to obtain a new system, which is then solved to check if the circuit is underconstrained. If the input variable causing the circuit to become underconstrained cannot be identified, the system of equations will be directly solved with Gröbner basis to verify the constrained situation of the circuit. The pseudo-code of checking polynomial circuits is given in Algorithm 4.
Algorithm 4 Algorithm for checking higher-order circuits
[…]
2:  $udcSols\leftarrow\textsc{GetUndeterminedCoeffSols}(K,U,E)$
3:  for all $sol\in udcSols$ do
4:   $E_{sub}\leftarrow\textsc{SubstitueEqs(E, sol)}$
5:   if Check $C(U,K,O,E_{sub})$ is underconstrained then
6:     return precisely underconstrained
7:   else
8:     continue
9:   end if
10:  end for
11:  $sols\leftarrow\textsc{SolveHigherSystemWith$F_{5}$}(C)$
12:  if IsEmpty(sols) then
13:   return algebraically overconstrained
14:  else if CheckUniqueness(sols, $O$) then
15:   return algebraically exact-constrained
16:  else if $\textsc{Size}(sols)>1$ then
17:   return precisely underconstrained
18:  else
19:   return unknown
20:  end if
21:end procedure